Let's try to use zoom
Amruthe reported about the weekly shift that was smooth.
Dayron remind the CMS request to have the gas line for charge measurements in H8.
As Luca reported that strating from sunday night it is expected an intervention on the electric system that will last all the week, Dayron will try to understand if it is not better to postpone the cms mearurement after this intervention.
In any case we should stop the system this sunday for one week.
Gianluca noted that EPDT chamber was at wrong voltage during aging runs and asked to increase the voltage.
Luca presnted the status of the 2022 Test beam plots (see indico)
There was a discussion about the current to use to correct the plateau curve for the drop across the electrodes. At moment we are using the current out of spill, but maybe that during the spill should be used. Luca will look at this.
Some of the cluster size measurement we are using for 2021 test beam plot seem to be strange for ECO2 EPDT chamber and for ECO2 and ECO3 of LHCb chamber. A check is required.